# Foundation Inference Models for Markov Jump Processes

### v2

@@ Acknowledgements
This research has been funded by the Federal Ministry of Education and Research of Germany and the state of North-Rhine Westphalia as part of the Lamarr-Institute for Machine Learning and Artificial Intelligence. César Ojeda is supported by Deutsche Forschungsgemeinschaft (DFG) - Project-ID 318763901 - SFB1294.
@@ A.1 Background on Markov Jump Processes in Continuous Time
Markov Jump Processes (MJPs) are stochastic models used to describe systems that transition between states at random times. These processes are characterized by the Markov property where the future state depends only on the current state, not on the sequence of events that preceded it.
@@ B.1 Prior Distributions and their Implementation
Distribution over rate matrices. Our data generation procedure starts by sampling the entries $f_{ij}$ of the intensity matrix from a beta distribution $f_{ij}\sim\text{Beta}(\rho_{f}=(\alpha,\beta))$, where $\alpha\in\{1,2\}$ and $\beta\in\{1,3,5,10\}$ are sampled uniformly. That is our prior $p(\rho_{f})$. The choices for $\alpha$ and $\beta$ were made heuristically to obtain reasonable distributions over the number of jumps (see Figure 8). We remark that we fixed this set of training distributions before evaluating the model on the evaluation sets, in order to prevent us from introducing a bias to the distribution hyperparameters by optimizing on the evaluation set. We set the adjacency matrix $\mathbf{A}=J$ with probability $\rho_{\text{\tiny fully\_connected}}=0.5$, where $J$ denotes the matrix for which all entries are $1$. For the remaining entries we apply a Bernoulli filter $\mathbf{A}=p(\rho_{\text{\tiny survival}})J$ that sets entries to 00 with probability $\rho_{\text{\tiny survival}}=0.5$. Our motivation for this is that it often happens in real world processes that the intensity matrices are sparse. We also remark that we only accept this masked-out matrix if the corresponding graph is connected, that is, if the system cannot get stuck in a single state.
We remark that while all entries of the intensity matrix during training were in the interval $[0,1]$, the model can still predict intensities beyond this interval (which we demonstrate on several evaluation sets), because we normalize the maximum time among the paths to be 1 and rescale the predicted intensities accordingly. Our approach for sampling intensity matrices resulted in a vast variety of different processes. The distribution of the number of jumps per trajectory is shown in figure 8. The distribution of relaxation times is shown in figure 9.
Distribution over initial conditions. We choose half of our initial distributions to be the stationary distribution of the MJP. The motivation for this is that it often happens that real life experiments produce very long observations of a system in equilibrium which then gets split into chunks in order to be processed. The second half of our initial distributions $\pi_{0}$ are randomly sampled from Dirichlet distributions $\pi_{0}\sim\text{Dir}(\rho_{0})$ where we heuristically choose $\rho_{0}=50$.
Distribution over observation grids. In practice, the exact jump times are not known. We therefore create observations of the current state of our system at a given time on a regular grid of 100 points. In order to make our model grid independent, we mask out grid points. Half of our grids are strided with $\text{strides}\in\{1,2,3,4\}$ and a Bernoulli filter with $\rho_{\text{survival}}\in\{1/4,1/2\}$ is applied to the other half in order to obtain irregular grids.
Distribution over noise process. Because real world data is often noisy we also add noise to the labels. If a state observation is selected to be mislabeled, the new label is randomly chosen from a uniform distribution over all states. We investigate two different configurations in this project, one with 1% label noise ($\rho_{y}=0.01$) and one with 10% label noise ($\rho_{y}=0.1$).
@@
Appendix C Model Architecture and Experimental Setup
@@ C.2 Experimental Setup
Training procedure: All models were trained on two A100 80Gb GPUs for approximately 500 epochs or approximately 2.5 days on average per model. Early stopping was employed as the stopping criterion. The models were trained by maximizing the likelihood. The implementation was done in Jax [^3] https://jax.readthedocs.io/en/latest/index.html. The code and trained models are also available online.
@@
How to use the model. The model takes the following inputs:
  • The observation grids ([num_paths, grid_size]): The normalized observation grids
  • The observation values ([num_paths, grid_size]): The observation values as an array of integers
  • The sequence lengths ([num_paths]): The number of points on the observation grids that the model should process (we used this during training to apply masks)
  • The time normalization factor: A float that scales the normalized times back to the original ones
The model then returns a $6\times 6$ intensity matrix (for systems with less than six states, one can just select the valid entries, although the model typically already predicts values close to zero for the non-occuring states). We recommend to use the model only inside its training range, that is with up to 300 paths and grids up to size 100.
Appendix D Ablation Study
@@
D.1 General Remarks about the Error Bars and Context Number
@@
D.2 Performance of the Model by varying its Architecture
The ablation study presented in Table 2 evaluates the impact of different model features on the performance by comparing various combinations of architectures and attention mechanisms with varying numbers of paths, and their corresponding RMSE values. The study examines models using a BiLSTM or Transformer, with and without self-attention and learnable query attention, across 1, 100, and 300 paths. The results indicate that increasing the number of paths consistently reduces RMSE (see section D.4 for more details), demonstrating the benefit of considering more paths during training. Specifically, using a BiLSTM with learnable query attention achieves an RMSE of $0.193\pm 0.031$ with a single path, significantly improving to $0.048\pm 0.011$ with 100 paths, and further to $0.0457\pm 0.0$ with 300 paths. Similarly, a Transformer with learnable query attention shows an RMSE of $0.196\pm 0.031$ for a single path, $0.049\pm 0.011$ for 100 paths, and $0.0458\pm 0.0$ for 300 paths. The inclusion of self-attention in the Transformer models slightly improves performance, with the best RMSE of $0.0459\pm 0.0$ achieved when both self-attention and learnable query attention are used with 300 paths. In this case since many of the processes contain one path it is beneficial to use the learnable query attention over the standard self-attention mechanism.
@@ D.2 Performance of the Model by varying its Architecture
Figure 10: Figure X: Impact of Hyperparameters on RMSE. The figure shows four line plots illustrating the effect of hyperparameters on model RMSE. The first plot shows RMSE increases with larger $\psi_{1}$ hidden sizes, being lowest at 256. The second plot indicates lower RMSE with a larger $\phi_{1}$ architecture size ([2x128]). The third plot shows minimal RMSE impact from $\phi_{2}$ architecture size. The fourth plot shows RMSE stability across different $\Omega_{1}$ hidden sizes, with slight variations based on $\psi_{1}$. This highlights the importance of tuning $\psi_{1}$ and $\phi_{1}$ for optimal performance.
@@
D.3 Performance of the Model with varying Number of States
@@
D.4 Performance of the Model with varying Number of Paths during Evaluation
@@
Table 5: Performance of FIM-MJP 1% given varying number of paths during the evaluation on the DFR dataset with regular grid. (LQ) denotes learnable-query-attention (see section C.1), (SA) denotes self-attention.
@@
Appendix E Additional Results
This section contains more of our results which did not fit into the main text. We begin this section by providing more details on the Hellinger distance which we used as a metric to assess the performance of our models. Afterwards, we provide more results and background on the ADP, ion channel and DFR datasets. Additionally, we introduce two two-state MJPs, given by the protein folding datasets (E.5) and the two-mode switching system (E.6), which we use to evaluate our models and to compare it against previous works.
E.1 Hellinger Distance
@@ E.4 DFR
E.5 Modeling Protein Folding through Bistable Dynamics
@@
We use the dataset of Seifner and Sánchez (2023), which was provided to us via private communication, and apply a Gaussian-Mixture-Model to classify the dataset into two states. The decision boundary of the classifier seems to be based on the absolute absolute value of the radius, namely the classifier seems to classify all states with a radius smaller than approximately 2 into the lower state (see figure 14).
@@
E.6 A Toy Two-Mode Switching System
@@
with parameters $\alpha_{1}=\alpha_{2}=1.5$, $\beta_{1}=-1$, and $\beta_{2}=1$. For a concise overview of the generation process, the reader is directed to (Köhs et al., 2021) for comprehensive details. We use the same dataset that was generated in (Seifner and Sánchez, 2023) using the code of (Köhs et al., 2021) which contains 256 paths of length 67 to evaluate our model. This dataset was provided to us via private communication. Our results are shown in table 16.

### v3

@@ Acknowledgements
This research has been funded by the Federal Ministry of Education and Research of Germany and the state of North-Rhine Westphalia as part of the Lamarr Institute for Machine Learning and Artificial Intelligence. Additionally, César Ojeda was supported by Deutsche Forschungsgemeinschaft (DFG) – Project-ID 318763901 – SFB1294.
We would like to thank Lukas Köhs for sharing the experimental ion channel data with us. The actual experiment was carried out by Kerri Kukovetz and Oliver Rauh while working in the lab of Gerhard Thiel of TU Darmstadt. Similarly, we would like to thank Nick Charron and Cecilia Clementi, from the Theoretical and Computational Biophysics group of the Freie Universität Berlin, for sharing the all-atom alanine dipeptide simulation data with us. The simulation was carried out by Christoph Wehmeyer while working in the research group of Frank Noé of the Freie Universität Berlin.
@@ A.1 Background on Markov Jump Processes in Continuous Time
Markov jump processes are stochastic models used to describe systems that transition between states at random times. These processes are characterized by the Markov property where the future state depends only on the current state, not on the sequence of events that preceded it.
@@ B.1 Prior Distributions and their Implementation
Distribution over rate matrices. Our data generation procedure starts by sampling the entries $f_{ij}$ of the intensity matrix from the following beta distributions
$p(f_{ij}|\rho_{f})=\text{Beta}(\rho_{f}=(\alpha,\beta)),\,\,\text{with}\,\,p(\alpha)=\text{Uniform}(\{1,2\})\\ \text{and}\,\,p(\beta)=\text{Uniform}(\{1,3,5,10\}).$ (14)
Both these discrete uniform distribution define the prior $p(\rho_{f})=p(\alpha)p(\beta)$.
The choices for $\alpha$ and $\beta$ were made heuristically, to obtain reasonable (i.e. varied) distributions over the number of jumps (see e.g. Figure 8). We remark that we fixed this set of training distributions before evaluating the model on the evaluation sets, in order to prevent us from introducing unwanted biases into the distribution hyperparameters by optimizing on the evaluation set.
Next we define the prior over the adjacency matrix as
[…]
Our motivation for this prior is that it often happens in real world processes that the intensity matrices are not fully connected. Let us remark, however, that we only accept the Erdös-Rényi sample if the corresponding graph is connected — that is, if the system cannot get stuck into a single state. Both these distributions implicitly define $p_{\text{\tiny rates}}(\mathbf{F}|\mathbf{A},\rho_{f})$, for $F_{ij}=a_{ij}f_{ij}$.
Remark on generalization beyond prior rate distribution. We remark that while all entries of the intensity matrix seen during training lie on the interval $[0,1]$, the model can still predict intensities outside this interval. We empirically demonstrated that this in indeed the case on the widely different target sets of the experimental section, in the main text. The reason behind this is that we normalize the maximum time among all input paths to be 1, and rescale the predicted intensities accordingly. Ultimately, what matters is the difference among the rates (and therefore among the observation times) within the target time series. Our approach for sampling intensity matrices resulted in a vast variety of different processes.
The distribution of the number of jumps per trajectory is shown in figure 8 and that of relaxation times is shown in figure 9.
@@
Distribution over initial conditions. We choose half of our initial distributions in our synthetic ensemble to be the stationary distribution of the MJP $p^{*}_{\text{\tiny MJP}}$. The motivation for this is that it often happens that real life experiments produce very long observations of a system in equilibrium. The second half of our initial distributions $\boldsymbol{\pi}_{0}$ are randomly sampled from a Dirichlet distribution $\text{Dir}(\rho_{0})$, where we heuristically choose $\rho_{0}=50$. In equations, we write
$p(\boldsymbol{\pi}_{0})=\frac{1}{2}p^{*}_{\text{\tiny MJP}}(\mathbf{F})+\frac{1}{2}\text{Dir}(\rho_{0}=50).$ (16)
Distribution over observation grids. In practice, the exact jump (i.e. transition) times are not known. We therefore first generate observations of the state of the system on a regular grid with a maximum of $L=100$ points. We then randomly mask out some observations from this fixed regular grid, in order to make the model grid independent. Half of our (subsampled) observation grids are chosen to be regular, i.e. they are strided with $\text{strides}\in\{1,2,3,4\}$. The other half are chosen to be irregular, through a Bernoulli filter (or mask) with $\rho_{\text{survival}}\in\{1/4,1/2\}$ applied to the base ($L=100$) grid.
Distribution over noise process. Because real world data is often noisy we also add noise to the labels. If a state observation is selected to be mislabeled, the new label is randomly chosen from a uniform distribution over all states. We investigate two different configurations in this project, one with 1% label noise ($\rho_{x}=0.01$) and one with 10% label noise ($\rho_{x}=0.1$).
@@
Appendix C How to use the Model: Inputs, Outputs and Rescalings
In this section, we give details about the inputs to and outputs of our pretrained recognition model. We also comment on the internal rescalings done by the model, in order to be able to infer MJPs from time series with observation times of any scale.
C.1 Input
The model takes as input three parameters:
  1. 1.
    The observation grids (shape: [num_paths $K$, grid_size $L$]): The observation times $\{\tau_{k1},\dots,\tau_{kl}\}_{k=1}^{K}$, padded to the maximum length $L$.
  2. 2.
    The observation values (shape: [num_paths $K$, grid_size $L$]): The noisy observation values $\{x^{\prime}_{k1},\dots,x^{\prime}_{kl}\}_{k=1}^{K}$ padded to the maximum length $L$. Note that these values are integers lying on the discrete set $\{0,1,\dots,C-1\}$.
  3. 3.
    The dimension $c$: The (a priori known) dimension of the process as an integer between 2 and $C$. If this dimension is unknown, the model returns a $C\times C$ rate matrix whose rank might (approximately) be smaller than $C$, which indicates a hidden state-space of size smaller than $C$.
We recommend users to use the model only within its training range. That is, with up to a maximum of $K=300$ paths, and grids up to a maximum of $L=100$ points.
C.2 Internal Rescaling
Internally, the model does the following:
  1. 1.
    It computes the maximum observation time:
    $\tau_{\text{\tiny max}}=\text{max}\{\tau_{k1},\dots,\tau_{kl}\}_{k=1}^{K}.$ (17)
  2. 2.
    It normalizes the observation times between 0 and 1:
    $\{\tau_{k1},\dots,\tau_{kl}\}_{k=1}^{K}\leftarrow\{\tau_{k1},\dots,\tau_{kl}\}_{k=1}^{K}/\tau_{\text{\tiny max}}.$ (18)
  3. 3.
    It computes the inter-event times $\Delta\tau_{ki}=\tau_{k,i+1}-\tau_{ki}$, for $k=1,\dots,K$.
  4. 4.
    It transforms the observation values to one-hot-encodings.
  5. 5.
    It predicts the (normalized) off-diagonal elements of the intensity matrix and variance matrix as well as the initial distribution (here we are working with the maximum supported dimension, that is $C$).
  6. 6.
    It rescales back the estimates to the original time scale:
[…]
    $\textrm{Var}\,\mathbf{\hat{F}}\leftarrow\textrm{Var}\,\mathbf{\hat{F}}/\tau_{\text{\tiny max}}.$ (20)
Note that, as we empirically demonstrated in the paper, this rescaling procedure allows us to work with real-world MJPs of arbitrary time scales. For example, the time scales for the switching ion channel dataset were more than 500 times smaller than the time scales in our training dataset.
C.3 Support for varying State Space Sizes
We now elaborate on how the model can deal with processes whose state spaces have sizes $c<C$.
We arranged all the target rate matrices $\mathbf{F}$ within our training dataset, for MJPs with state spaces of size $c<C$, to be the leftmost block diagonal $c\times c$ matrix within a $C\times C$ matrix of zeros, so that the redundant matrix elements are always zero. As can be read from equation (6) of the main text, we train FIM to predict zeros for those redundant matrix elements.
In practice, however, our trained FIM does not exactly predict zeros for those redundant matrix elements. In our experiments, the user knows a priori the number of states $c$ of the hidden process, so we explicitly set the redundant matrix elements to zero, and only then compute the corrected diagonal (i.e. the normalization) of the output rate matrix.
We afterwards select the $c\times c$ entries of the predicted intensity matrix and variance matrix as well as the first $c$ entries of the predicted initial distribution.
We refer the reader to our library for additional details.
C.4 Output
The output of the model consists of three parameters:
  1. 1.
    The intensity matrix $\mathbf{\hat{F}}$ (shape: [$c$, $c$]).
  2. 2.
    The variance matrix $\textrm{Var}\,\mathbf{\hat{F}}$ (shape: [$c$,$c$]).
  3. 3.
    The initial distribution $\pi_{0}$ (shape [$c$]).
Appendix D Model Architecture and Experimental Setup
@@ D.2 Experimental Setup
Training procedure: All models were trained on two A100 80Gb GPUs for approximately 500 epochs or approximately 2.5 days on average per model. Early stopping was employed as the stopping criterion. The models were trained by maximizing the likelihood.
@@
Pretrained models: Our pretrained models are also available online [^5] https://github.com/cvejoski/OpenFIM.
Appendix E Ablation Studies
@@
E.1 General Remarks about the Error Bars and Context Number
@@
E.2 Performance of the Model by varying its Architecture
The ablation study presented in Table 2 evaluates the impact of different model features on the performance by comparing various combinations of architectures and attention mechanisms with varying numbers of paths, and their corresponding RMSE values. The study examines models using a BiLSTM or Transformer, with and without self-attention and learnable query attention, across 1, 100, and 300 paths. The results indicate that increasing the number of paths consistently reduces RMSE (see section E.4 for more details), demonstrating the benefit of considering more paths during training. Specifically, using a BiLSTM with learnable query attention achieves an RMSE of $0.193\pm 0.031$ with a single path, significantly improving to $0.048\pm 0.011$ with 100 paths, and further to $0.0457\pm 0.0$ with 300 paths. Similarly, a Transformer with learnable query attention shows an RMSE of $0.196\pm 0.031$ for a single path, $0.049\pm 0.011$ for 100 paths, and $0.0458\pm 0.0$ for 300 paths. The inclusion of self-attention in the Transformer models slightly improves performance, with the best RMSE of $0.0459\pm 0.0$ achieved when both self-attention and learnable query attention are used with 300 paths. In this case since many of the processes contain one path it is beneficial to use the learnable query attention over the standard self-attention mechanism.
@@ E.2 Performance of the Model by varying its Architecture
Figure 10: Impact of Hyperparameters on RMSE. The figure shows four line plots illustrating the effect of hyperparameters on model RMSE. The first plot shows RMSE increases with larger $\psi_{1}$ hidden sizes, being lowest at 256. The second plot indicates lower RMSE with a larger $\phi_{1}$ architecture size ([2x128]). The third plot shows minimal RMSE impact from $\phi_{2}$ architecture size. The fourth plot shows RMSE stability across different $\Omega_{1}$ hidden sizes, with slight variations based on $\psi_{1}$. This highlights the importance of tuning $\psi_{1}$ and $\phi_{1}$ for optimal performance.
@@
E.3 Performance of the Model with varying Number of States
@@
E.4 Performance of the Model with varying Number of Paths during Evaluation
@@
Table 5: Performance of FIM-MJP 1% given varying number of paths during the evaluation on the DFR dataset with regular grid. (LQ) denotes learnable-query-attention (see section D.1), (SA) denotes self-attention.
@@
Appendix F Additional Results
This section contains more of our results which did not fit into the main text. We begin this section by providing more details on the Hellinger distance which we used as a metric to assess the performance of our models. Afterwards, we provide more results and background on the ADP, ion channel and DFR datasets. Additionally, we introduce two two-state MJPs, given by the protein folding datasets (F.5) and the two-mode switching system (F.6), which we use to evaluate our models and to compare it against previous works.
F.1 Hellinger Distance
@@ F.4 Discrete Flashing Ratchet
F.5 Modeling Protein Folding through Bistable Dynamics
@@
We use the dataset of Seifner and Sánchez (2023) and apply a Gaussian-Mixture-Model to classify the dataset into two states. The decision boundary of the classifier seems to be based on the absolute absolute value of the radius, namely the classifier seems to classify all states with a radius smaller than approximately 2 into the lower state (see figure 14).
@@
F.6 A Toy Two-Mode Switching System
@@
with parameters $\alpha_{1}=\alpha_{2}=1.5$, $\beta_{1}=-1$, and $\beta_{2}=1$. For a concise overview of the generation process, the reader is directed to (Köhs et al., 2021) for comprehensive details. We use the same dataset that was generated in (Seifner and Sánchez, 2023) using the code of (Köhs et al., 2021) which contains 256 paths of length 67 to evaluate our model. Our results are shown in table 17.